OpenStudy (anonymous):

For the given cost function "C(x)=19600+300x+x^2" find: a) The production level that will minimize the average cost b) The minimal average cost Any help is appreciated. TIA

OpenStudy (anonymous):

u knw hw to differentiate the function?

OpenStudy (anonymous):

Yes I do. C'(x) = 2(x+150) I believe

OpenStudy (anonymous):

\[C'(x)=300 + 2x\]

OpenStudy (anonymous):

it is indeed. on differentiating it for a second time, u get just 2, which is greater than zero which holds true for minima. so equate the first derivative to zero & find x \[2(x+150)=0\]

OpenStudy (anonymous):

x = -150 in that case

OpenStudy (anonymous):

which makes x = -150. put that in C(x) function. that will give u the min cost. as for production level, m not sure about that to be honest.

OpenStudy (anonymous):

Ah, no worries, thank you for your help!

OpenStudy (anonymous):

welcome... but did u understand the production level question anyway?

OpenStudy (anonymous):

Actually, I just tried it, and it appears to be incorrect. I plugged "-150" into the C(x) function to get "-2900".

OpenStudy (anonymous):

yes. seems theres something wrong. hold on!

OpenStudy (anonymous):

okay, m pretty sure that the method followed is correct. minimizing the function does give -150. ur question seems to have some fault. re check it

OpenStudy (anonymous):

I've looked over my question and it is worded exactly as in the original post :/

OpenStudy (anonymous):

then the question itself is wrong. i just checked with wolframalpha. computers cant be wrong :/ http://www.wolframalpha.com/input/?i=minimize+19600+%2B+300x+%2Bx%5E2

OpenStudy (anonymous):

I suspect it has something to do with the average cost? Because right now, we're only deriving the cost function and not working with the average cost...

OpenStudy (anonymous):

When you are minimizing average cost, you need to divide the cost function by x then differentiates and set it equal to zero to fine the minimal

OpenStudy (anonymous):

now if u'd only mind explaining what u just said? no pun intended.

OpenStudy (anonymous):

"19600+300x+x^2" / "x" = "(19600/x)+300+x" Derivative of: "(19600/x)+300+x" is "1-(19600/x^2)" 1-(19600/x^2)?

OpenStudy (anonymous):

its by definition of average you add up all the numbers and divide by the amount you are only given the cost function not the average cost function

OpenStudy (anonymous):

yea that is the correct derivative

OpenStudy (anonymous):

part a is the optimal x value and part b is the optimal y value

OpenStudy (anonymous):

solutions will be -1/140 & +1/140

OpenStudy (anonymous):

And to get the desired result of 0, we just find the square-root of 19600, so that the 1 can subtract into 1 to get 0? So the answer should be "140"?

OpenStudy (anonymous):

& on finding the second derivative: \[C''(x)=-39200x\]

OpenStudy (anonymous):

so the value of x that will be used in the original function(the one that has been divided by x) is -1/140

OpenStudy (anonymous):

the answer i got is 140 for the optimal x value

OpenStudy (anonymous):

it doesnt really matter in this case since there is a 'square' but u hav to keep that in mind

OpenStudy (anonymous):

no wait, i screwed up again :P

OpenStudy (anonymous):

140 is indeed the value to be used! @jayz657

OpenStudy (anonymous):

While we know that 140 is the optimal x value, I'm uncertain as to how we can find out the minimal average cost.

OpenStudy (anonymous):

\[C''(x)=39200/x^3\]

OpenStudy (anonymous):

the minimal average cost is when you just plug in 140 for the x value

OpenStudy (anonymous):

Do we just plug 140 into the cost function?

OpenStudy (anonymous):

yea thats right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

answer is 0 btw :3

OpenStudy (anonymous):

you need to do C(140)

OpenStudy (anonymous):

Average cost function that is. The answer is 580 I believe.

OpenStudy (anonymous):

your answer should be bigger than 19,200

OpenStudy (anonymous):

19600 sorry

OpenStudy (anonymous):

so u plugged 140 in \[19600 + 300x + x^2 \] or \[19600/x +300 +x\]

OpenStudy (anonymous):

crap sorry i messed up your right its 580

OpenStudy (anonymous):

its the second equation

OpenStudy (anonymous):

We are looking for: b) The minimal average cost So I plugged "140" into the average cost function and got "580"

OpenStudy (anonymous):

Ah, glad we're on the right page. Thank you for your help today folks, it's appreciated. :)

OpenStudy (anonymous):

your welcome glad to help

OpenStudy (anonymous):

@jayz657 thanks from my side too ^_^ !